Consider the following reaction:
2HBr(g)→H2(g)+Br2(g)

Part B

In the first 22.0 s of this reaction, the concentration of HBr dropped from 0.570 M to 0.493 M . Calculate the average rate of the reaction in this time interval.

Express your answer using two significant figures.

Rate

Rate

=

nothing

  M⋅s−1  
Answer #1

2HBr(g)→H2(g)+Br2(g)

rate = -1/2 [HBr] / t

[HBr] = 0.493 - 0.570 = -0.077

t = 22.0 s

rate = -1/2 [-0.077] / 22.0

rate = 3.5 x 10-3 M.s-1